OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form.

OpenStudy (anonymous):

y=mx+b where m is slope. here: y = 3x + b. Because line passes through 2,5 you get b by using these values for x and y.

OpenStudy (anonymous):

5 = 3*2 + b b = -1 So line is: y = 3x -1

OpenStudy (anonymous):

??

OpenStudy (anonymous):

its wrong the answer that you gave me is not on the options of the test thanku though!

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

A.y + 5 = 3(x + 2) B.y – 5 = 3(x – 2) C.y + 8 = 3(x + 5) D.y – 2 = 3(x – 5)

OpenStudy (anonymous):

B. It is the same equation expressed differently.

OpenStudy (anonymous):

y-5 = 3(x-2) y - 5 = 3x - 6 y = 3x - 6 + 5 y = 3x - 1

OpenStudy (anonymous):

Medal?